# Tight inapproximability of max-LINSAT and implications for decoded quantum interferometry

### 04 Mar 2026 v1

@@ Abstract
We establish tight inapproximability bounds for max-LINSAT, the problem of maximizing the number of satisfied linear constraints over the finite field $\mathbb{F}_{q}$, where each constraint accepts $r$ values. Specifically, we prove by a direct reduction from Håstad’s theorem that no polynomial-time algorithm can exceed the random-assignment ratio $r/q$ by any constant, assuming $\mathsf{P}\neq\mathsf{NP}$. This threshold coincides with the $\ell/m\to 0$ limit of the semicircle law governing decoded quantum interferometry (DQI), where $\ell$ is the decoding radius of the underlying code: as the decodable structure vanishes, DQI’s approximation ratio degrades to exactly the worst-case bound established by our result. Together, these observations delineate the boundary between worst-case hardness and potential quantum advantage, showing that any algorithm surpassing $r/q$ must exploit algebraic structure specific to the instance.
@@ 1 Introduction
DQI targets a generalized linear constraint satisfaction problem over finite fields, commonly referred to as max-LINSAT. The flagship application is optimal polynomial intersection (OPI) problem, a structured max-LINSAT instance whose constraint matrix has Vandermonde structure arising from Reed–Solomon codes that are efficiently decodable up to their designed error radius. For OPI, in a certain parameter regime, DQI achieves a super-polynomial speed-up over known classical algorithms.
A growing body of follow-up work has developed around DQI [14, 32, 37, 10, 9, 39, 30, 2, 34, 31, 19, 28, 40, 23, 26, 15, 38, 11]. Most relevant to our setting, the DQI advantage for OPI has been strengthened via soft decoding of Reed–Solomon codes [14, 15] and extended to algebraic geometry codes via a generalization called HOPI [19]. At the same time, several works have exposed DQI’s limitations [31, 2, 30].
A prerequisite for precisely interpreting the results surrounding DQI is a clear understanding of the approximability limits of the underlying optimization problem. For many well-studied constraint satisfaction problems, tight inapproximability thresholds are known, often matching the performance of a random assignment. For max-LINSAT in the formulation considered by DQI, which allows linear constraints with arbitrary acceptance sets over a finite field, no such worst-case guarantees have previously been established. It is, therefore, unclear whether the approximation ratios achieved by quantum or classical heuristics on generic instances are close to being optimal, or whether substantially better algorithms might exist.
@@
Our main result shows that this random baseline is provably optimal in the worst case. Specifically, for every finite field $\mathbb{F}_{q}$ of order $q$, every $1\leq r\leq q-1$, and every $\varepsilon>0$, it is $\mathsf{NP}$-hard to distinguish between instances in which almost all constraints are satisfiable and instances in which no assignment satisfies more than a $(r/q+\varepsilon)$-fraction of constraints. Assuming $\mathsf{P}\neq\mathsf{NP}$, no polynomial-time classical algorithm can approximate max-LINSAT beyond the random-assignment threshold, and likewise no polynomial-time quantum algorithm can do so unless $\mathsf{NP}\subseteq\mathsf{BQP}$. From the perspective of quantum optimization, this provides a complexity-theoretic benchmark. The threshold $r/q$ is hence a hard wall for worst-case max-LINSAT(q,r)(q,r): no efficient algorithm can consistently exceed it on arbitrary instances. This does not conflict with DQI’s demonstrated advantage on OPI, where the Vandermonde structure of classical Reed–Solomon codes enables approximation ratios far above the random baseline. Rather, our result delineates the boundary: any algorithm that beats $r/q$ on a max-LINSAT instance must exploit structure specific to that instance. The potential for quantum advantage in this setting lies in leveraging algebraic structure rather than overcoming worst-case complexity barriers.
The proof builds on and generalizes an inapproximability result of Håstad for max-E33-LIN over finite Abelian groups, which establishes tight hardness for the case of single-valued acceptance sets [21]. We show how this hardness extends to max-LINSAT instances with uniform acceptance sets of size $r$, yielding a matching hardness threshold for the general problem. The resulting inapproximability bounds are tight.
@@ 2 Preliminaries
$\sum_{j\in S_{i}}c_{i,j}x_{j}$ $=b_{i},\qquad i=1,\dots,m,$ (1)
where for each $i$, $S_{i}\subseteq\{1,\dots,n\}$ satisfies $|S_{i}|=k$, $c_{i,j}\in\mathbb{F}_{q}\setminus\{0\}$ for all $j\in S_{i}$, and $b_{i}\in\mathbb{F}_{q}$. The goal is to find an assignment $\boldsymbol{x}\in\mathbb{F}_{q}^{n}$ that satisfies the maximum number of equations.
To establish worst-case limits on the approximation for max-LINSAT, we rely on a foundational result of Håstad [21, Theorem 5.9], which proves strong inapproximability for max-E33-LIN-$\Gamma$ over finite Abelian groups. Håstad’s result builds on the celebrated PCP theorem [4, 3], which characterizes $\mathsf{NP}$ in terms of probabilistically checkable proofs and underlies most modern inapproximability results. This result provides the base case from which we extend hardness to max-LINSAT instances with uniform acceptance sets of arbitrary size.
@@
In particular, setting $\Gamma=(\mathbb{F}_{q},+)$ for a finite field $\mathbb{F}_{q}$ yields the $(1-\varepsilon,\;1/q+\varepsilon)$-hardness of max-E33-LIN-$q$. Since each coefficient $a_{i}$ is invertible in $\mathbb{F}_{q}$, the substitution $y_{i}=a_{i}x_{i}$ reduces any instance with coefficients to the coefficient-free form (i.e., all coefficients equal to one) covered by the theorem.
@@
The inapproximability of max-LINSAT(q,r)(q,r) can also be situated within the broader theory of constraint satisfaction problems. Under the unique games conjecture (UGC) [27], Ref. [36] showed that for every constraint satisfaction problem (CSP), the optimal polynomial-time approximation ratio equals the integrality gap of a canonical semi-definite programming (SDP) relaxation; moreover, this ratio is achieved by an appropriate rounding of the SDP solution. More concretely, Ref. [5] has proven that any predicate whose satisfying assignments support a balanced pairwise-independent distribution is approximation resistant under the UGC. The max-LINSAT(q,r)(q,r) predicate for arity $k=3$, i.e., checking whether $x_{i_{1}}+x_{i_{2}}+x_{i_{3}}$ falls in an $r$-element subset $S\subseteq\mathbb{F}_{q}$, satisfies this criterion (as does the analogous predicate for any fixed $k\geq 3$), so UGC-hardness of beating $r/q$ follows immediately from their result. Our Theorem 5 establishes the same conclusion under the weaker assumption $\mathsf{P}\neq\mathsf{NP}$, via a direct reduction from Håstad’s theorem [21]. This is notable because the UGC remains open. In a different direction, Ref. [16] proved unconditional ($\mathsf{P}\neq\mathsf{NP}$) approximation resistance for predicates whose satisfying assignments contain a pairwise-independent subgroup. For prime $q$, however, the only additive subgroups of $\mathbb{F}_{q}$ are $\{0\}$ and $\mathbb{F}_{q}$ itself, so acceptance sets of size $1\leq r\leq q-1$ are never subgroups, and this criterion does not apply to our setting. For prime power $q$, nontrivial additive subgroups do exist, but the criterion still does not cover generic acceptance sets of arbitrary size $r$. Ref. [18] extended Håstad’s max-E33-LIN hardness from Abelian to all finite groups, and Ref. [13] recently transferred these inapproximability results to the setting of promise constraint satisfaction problems. These results may prove useful for establishing inapproximability bounds for potential DQI variants over more general finite groups.
@@ 3 Results
The case $r=1$ follows directly from Theorem 4, since every instance of max-E33-LIN-$q$ (with coefficients) is syntactically an instance of max-LINSAT(q,1)(q,1): each equation $\sum_{j\in S_{i}}c_{i,j}x_{j}=b_{i}$ defines a constraint with $F_{i}={b_{i}}$.
@@
Since $(q-r)/(q-1)\leq 1$, this is at most $r/q+\varepsilon$. Therefore $\mathrm{OPT}(\phi^{\prime})\leq(r/q+\varepsilon)\,m^{\prime}$, completing the soundness analysis.
@@ 4 Discussion
A key observation is that DQI’s quantum advantage claims are not for arbitrary max-LINSAT instances, but for algebraically structured problems, most notably optimal polynomial intersection (OPI), where the constraint matrix has Vandermonde structure arising from classical Reed–Solomon codes [25]. For OPI with parameters $n/q=r/q=1/2$, DQI achieves an approximation ratio of approximately $0.933$ via the semicircle law, while the best known classical algorithm (Prange’s information set decoding [35]) achieves only ${\sim}0.55$ [25]. These insights have recently been extended to algebraic geometry codes in Ref. [19], where a problem called HOPI is studied, which generalizes OPI. Therefore, DQI targets structured instances, not worst-case max-LINSAT.
Our worst-case inapproximability result is entirely consistent with these claims: Theorem 5 applies to arbitrary max-LINSAT(q,r)(q,r) instances, whereas OPI instances possess algebraic structure that DQI exploits via coherent implementation of the Berlekamp–Massey decoding algorithm. The hardness reduction from max-E33-LIN-$q$ produces instances with no such exploitable structure.
In DQI, the optimization problem is mapped to a decoding task for a classical error-correcting code. A central parameter governing the algorithm’s performance is the decoding radius $\ell$, defined as the maximum number of errors that an efficient decoder can correct. If a max-LINSAT instance with $m$ constraints arises from a code with decoding radius $\ell$, then DQI coherently maps assignments within Hamming distance $\ell$ to codewords. The resulting approximation ratio is given by the semicircle law [25],
@@
Figure 1: Landscape of approximability for $\mathrm{max\text{-}LINSAT}(q,r)$. The x-axis parametrizes the decodable structure $\ell/m$, where $\ell$ is the decoding radius of the underlying error-correcting code and $m$ is the number of constraints. The shaded region, spanning approximation ratios from $r/q$ to 11 across all values of $\ell/m$, is $\mathsf{NP}$-hard to achieve on worst-case instances (Theorem 5). The pink curve depicts DQI’s approximation ratio as given by the semicircle law [25], which surpasses the random-assignment bound $r/q$ whenever $\ell/m>0$ and saturates at $\alpha_{\mathrm{DQI}}=1$ for $\ell/m\geq 1-r/q$. In the limit $\ell/m\to 0$, DQI’s performance degrades to exactly $r/q$, matching the worst-case barrier.
This observation is complemented by Theorem 5, which shows that on worst-case max-LINSAT(q,r)(q,r) instances, no polynomial-time algorithm can exceed the $r/q$ threshold (assuming $\mathrm{P}\neq\mathrm{NP}$). While these two statements operate in different regimes – the semicircle law describes DQI’s performance on code-derived instances as a function of the decoding radius, and Theorem 5 applies to adversarially constructed instances with no exploitable structure – they point to the same conclusion from different directions: $r/q$ is the natural worst-case approximation barrier for max-LINSAT, and surpassing it requires both an algorithm capable of exploiting algebraic structure and an instance that possesses such structure. The codes for which DQI demonstrates an advantage over known classical algorithms, namely Reed–Solomon codes in OPI [25] and algebraic geometry codes in HOPI [19], satisfy both conditions.
@@
Recent work has further clarified that DQI’s advantage, where it exists, depends critically on problem structure: Ref. [31] shows that for MaxCut instances where DQI achieves any nontrivial approximation guarantee, the optimal solution can be found exactly in classical polynomial time. DQI consequently provides no advantage for MaxCut. For unstructured max-$k$-XOR-SAT instances, Ref. [2] proves that DQI is obstructed by the overlap gap property, which is a topological barrier in the near-optimal solution space. Classical approximate message passing outperforms DQI on such instances. From the perspective of computational complexity, Ref. [30] places DQI in $\mathsf{BPP}^{\mathsf{NP}}$, thereby ruling out immediate quantum advantage or “supremacy” arguments of the type as they are used in quantum random sampling [20, 8]. There, it is $\mathsf{P}-$hard to compute individual output probabilities, and an efficient classical sampling algorithm up to a constant error in the total variation distance would lead to a collapse of the polynomial hierarchy to the third level. In contrast, DQI is already placed in a low level of the polynomial hierarchy. The hardness of DQI, if any, arises from locating elements in an exponentially large hidden subset without group structure and not from the hardness of simulating the quantum circuit itself, as here, in contrast to the situation in quantum random sampling, individual probabilities of outcomes can be efficiently computed.
Together with Theorem 5, these results rigorously establish that DQI is not a general-purpose quantum optimizer. Several natural questions remain:
  • Inapproximability for OPI and HOPI: Can inapproximability results be extended to settings studied in OPI [25] and HOPI [19]? The Vandermonde structure of OPI instances places them in a restricted subclass of max-LINSAT for which generic PCP reductions do not directly apply, and it remains open whether this structure makes the problem classically easier or merely differently hard. A positive answer would strengthen DQI’s advantage claim; a negative answer (efficient classical algorithms for OPI) would “dequantize” DQI.
@@
  • Quadratic constraints: Ref. [23] extends DQI to quadratic constraints over finite fields (max-QUADSAT). For $q=2$ with singleton acceptance sets, Håstad [22] establishes tight inapproximability at the $2^{-d}$ random-assignment threshold for degree-$d$ equations, giving $1/4$ for quadratic constraints, which often appear in practically relevant optimization problems. Extending this to general $\mathbb{F}_{q}$ and acceptance sets of size larger than one remains open: the proof of Ref. [22] relies on $\mathbb{F}_{2}$-specific structure (multilinearity and Reed–Muller minimum weight characterizations) that does not directly generalize.
@@
Our Theorem 5 provides a complexity-theoretic benchmark for max-LINSAT, establishing $r/q$ as the provable worst-case approximation limit and confirming that any improvement must arise from exploitable algebraic structure. It is our hope that this result contributes to clarifying the landscape of quantum advantage for optimization problems of practical relevance [17, 24, 1, 6].

### 22 Mar 2026 v2

@@ Abstract
We establish tight inapproximability bounds for max-LINSAT, the problem of maximizing the number of satisfied linear constraints over the finite field $\mathbb{F}_{q}$, where each constraint accepts $r$ values. Specifically, we prove by a direct reduction from Håstad’s theorem that no polynomial-time algorithm can exceed the random-assignment ratio $r/q$ by any constant, assuming $\mathsf{P}\neq\mathsf{NP}$. This threshold coincides with the $\ell/m\to 0$ limit of the semicircle law governing decoded quantum interferometry (DQI), where $\ell$ is the decoding radius of the underlying code. Together, these observations delineate the boundary between worst-case hardness and potential quantum advantage, showing that any algorithm surpassing $r/q$ must exploit instance structure beyond what is present in the hard instances produced by PCP reductions.
@@ 1 Introduction
DQI targets a generalized linear constraint satisfaction problem over finite fields, commonly referred to as max-LINSAT. The flagship application is the optimal polynomial intersection (OPI) problem, a structured max-LINSAT instance whose constraint matrix has Vandermonde structure arising from Reed–Solomon codes that are efficiently decodable up to their designed error radius. For OPI, in a certain parameter regime, DQI achieves a super-polynomial speed-up over known classical algorithms.
A growing body of follow-up work has developed around DQI [14, 31, 36, 10, 9, 38, 29, 2, 33, 30, 19, 27, 39, 25, 15, 37, 11]. Most relevant to our setting, the DQI advantage for OPI has been strengthened via soft decoding of Reed–Solomon codes [14, 15] and extended to algebraic geometry codes via a generalization called HOPI [19]. At the same time, several works have exposed DQI’s limitations [30, 2, 29].
A prerequisite for precisely interpreting the results surrounding DQI is a clear understanding of the approximability limits of the underlying optimization problem. For many well-studied constraint satisfaction problems, tight inapproximability thresholds are known, often matching the performance of a random assignment. For max-LINSAT in the formulation considered by DQI, which allows linear constraints with arbitrary acceptance sets over a finite field, no such worst-case guarantees have previously been established. Without such worst-case bounds, it is unclear whether the approximation ratios achieved by quantum or classical heuristics are close to being optimal, or whether substantially better polynomial-time algorithms might exist.
@@
Our main result shows that this random baseline is provably optimal in the worst case. Specifically, for every finite field $\mathbb{F}_{q}$ of order $q$, every $1\leq r\leq q-1$, and every $\varepsilon>0$, it is $\mathsf{NP}$-hard to distinguish between instances in which almost all constraints are satisfiable and instances in which no assignment satisfies more than a $(r/q+\varepsilon)$-fraction of constraints. Assuming $\mathsf{P}\neq\mathsf{NP}$, no polynomial-time classical algorithm can approximate max-LINSAT consistently beyond the random-assignment threshold, and likewise no polynomial-time quantum algorithm can do so unless $\mathsf{NP}\subseteq\mathsf{BQP}$. From the perspective of quantum optimization, this provides a complexity-theoretic benchmark. The threshold $r/q$ is hence a hard wall for worst-case max-LINSAT(q,r)(q,r): no efficient algorithm can consistently exceed it on arbitrary instances. This does not conflict with DQI’s demonstrated advantage on OPI, which operates in a qualitatively different regime with constant $n/m$, and instances derived from classical Reed–Solomon codes that exhibit specific Vandermonde structure. For such instances, both DQI and classical algorithms achieve approximation ratios exceeding the random assignment baseline, whereas the hard instances produced by PCP-reductions have $n/m=o(1)$. Rather, our result delineates a boundary: while structured instances may admit algorithms that beat the random-assignment threshold, no polynomial-time algorithm can do so on arbitrary max-LINSAT instances under reasonable complexity-theoretic assumptions.
Our proof builds on and generalizes an inapproximability result of Håstad for max-E33-LIN over finite Abelian groups, which establishes tight hardness for the case of single-valued acceptance sets [22]. We show how this hardness extends to max-LINSAT instances with uniform acceptance sets of size $r$, yielding a matching hardness threshold for the general problem. The resulting inapproximability bounds are tight.
@@ 2 Preliminaries
$\sum_{j\in S_{i}}x_{j}$ $=b_{i},\qquad i=1,\dots,m,$ (1)
where for each $i$, $S_{i}\subseteq\{1,\dots,n\}$ satisfies $|S_{i}|=k$, and $b_{i}\in\mathbb{F}_{q}$. The goal is to find an assignment $\boldsymbol{x}\in\mathbb{F}_{q}^{n}$ that satisfies the maximum number of equations.
To establish worst-case limits on the approximation for max-LINSAT, we rely on a foundational result of Håstad [22, Theorem 5.9], which proves strong inapproximability for max-E33-LIN-$\Gamma$ over finite Abelian groups. Håstad’s result builds on the celebrated PCP theorem [4, 3], which characterizes $\mathsf{NP}$ in terms of probabilistically checkable proofs and underlies most modern inapproximability results. This result provides the base case from which we extend hardness to max-LINSAT instances with uniform acceptance sets of arbitrary size.
@@
In particular, setting $\Gamma=(\mathbb{F}_{q},+)$ for a finite field $\mathbb{F}_{q}$ yields the $(1-\varepsilon,\;1/q+\varepsilon)$-hardness of max-E33-LIN-$q$.
@@
The inapproximability of max-LINSAT(q,r)(q,r) can also be situated within the broader theory of constraint satisfaction problems. Under the unique games conjecture (UGC) [26], Ref. [35] showed that for every constraint satisfaction problem (CSP), the optimal polynomial-time approximation ratio equals the integrality gap of a canonical semi-definite programming (SDP) relaxation; moreover, this ratio is achieved by an appropriate rounding of the SDP solution. More concretely, Ref. [5] has proven that any predicate whose satisfying assignments support a balanced pairwise-independent distribution is approximation resistant under the UGC. The max-LINSAT(q,r)(q,r) predicate for arity $k=3$ and coefficients being one, i.e., checking whether $x_{i_{1}}+x_{i_{2}}+x_{i_{3}}$ falls in an $r$-element subset $S\subseteq\mathbb{F}_{q}$, satisfies this criterion (as does the analogous predicate for any fixed $k\geq 3$), so UGC-hardness of beating $r/q$ follows immediately from their result. Our Theorem 5 establishes the same conclusion under the weaker assumption $\mathsf{P}\neq\mathsf{NP}$, via a direct reduction from Håstad’s theorem [22]. This is notable because the UGC remains disputed. In a different direction, Ref. [16] proved unconditional ($\mathsf{P}\neq\mathsf{NP}$) approximation resistance for predicates whose satisfying assignments contain a pairwise-independent subgroup. For prime $q$, however, the only additive subgroups of $\mathbb{F}_{q}$ are $\{0\}$ and $\mathbb{F}_{q}$ itself, so acceptance sets of size $1\leq r\leq q-1$ are never subgroups, and this criterion does not apply to our setting. For prime power $q$, nontrivial additive subgroups do exist, but the criterion still does not cover generic acceptance sets of arbitrary size $r$. Ref. [18] extended Håstad’s max-E33-LIN hardness from Abelian to all finite groups, and Ref. [13] recently transferred these inapproximability results to the setting of promise constraint satisfaction problems. These results may prove useful for establishing inapproximability bounds for potential DQI variants over more general finite groups.
@@ 3 Results
The case $r=1$ follows directly from Theorem 4, since every instance of max-E33-LIN-$q$ is syntactically an instance of max-LINSAT(q,1)(q,1) with coefficients $B_{i,j}$ being one if and only if the variable is involved: each equation $\sum_{j\in S_{i}}x_{j}=b_{i}$ defines a constraint with $F_{i}=\{b_{i}\}$.
@@
Since $(q-r)/(q-1)\leq 1$, this is at most $r/q+\varepsilon$. Therefore, $\mathrm{OPT}(\phi^{\prime})\leq(r/q+\varepsilon)\,m^{\prime}$, completing the soundness analysis. ∎
@@
Remark 7 (Coefficient-free hard instances).
The hard instances produced by the reduction have a particularly simple form: the constraint matrix $B$ has entries in $\{0,1\}$ only, with exactly three ones per row. This follows from the fact that Håstad’s construction yields coefficient-free equations $x_{i_{1}}+x_{i_{2}}+x_{i_{3}}=b_{i}$. The inapproximability of max-LINSAT(q,r)(q,r) therefore already holds for this restricted class of instances, which is a stronger statement than hardness for general constraint matrices.
@@ 4 Discussion
A key observation is that DQI’s quantum advantage claims are not for arbitrary max-LINSAT instances, but for structured problems, most notably optimal polynomial intersection (OPI), with constant $n/m$ and a constraint matrix that has Vandermonde structure arising from classical Reed–Solomon codes, which DQI exploits via coherent implementation of the Berlekamp–Massey decoding algorithm [24]. For OPI with parameters $n/q=r/q=1/2$, DQI achieves an approximation ratio of approximately $0.933$ via the semicircle law, while the best known classical algorithm (Prange’s information set decoding [34]) achieves only ${\sim}0.55$ [24]. These insights have recently been extended to algebraic geometry codes in Ref. [19], where a problem called HOPI is studied, which generalizes OPI. Therefore, DQI targets structured instances, not worst-case max-LINSAT.
Our worst-case inapproximability result is entirely consistent with these claims: Theorem 5 applies to worst-case max-LINSAT(q,r)(q,r) instances, whereas OPI operates in the structured regime described above. The hardness reduction from max-E33-LIN-$q$ produces instances with $n/m=o(1)$ and no such exploitable structure.
In DQI, the optimization problem is mapped to a decoding task for a classical error-correcting code. A central parameter governing the algorithm’s performance is the decoding radius $\ell$, defined as the maximum number of errors that an efficient decoder can correct. If a max-LINSAT instance with $m$ constraints gives rise to a dual code $C^{\perp}$ that can be decoded from up to $\ell$ errors, then DQI exploits this decodability to prepare a degree-$\ell$ polynomial enhancement of the objective function in superposition. The resulting approximation ratio is given by the semicircle law [24],
@@
Figure 1: Landscape of approximability for $\mathrm{max\text{-}LINSAT}(q,r)$. The x-axis parametrizes the decodable structure $\ell/m$, where $\ell$ is the decoding radius of the underlying error-correcting code and $m$ is the number of constraints. For worst-case instances, by virtue of Theorem 5, approximation ratios within the blue-shaded region are unattainable. Importantly, this hardness statement does not depend on the parameter $\ell/m$: the x-axis is only meaningful for structured instances arising from decoding problems (such as those targeted by DQI). The shading therefore illustrates a worst-case approximation barrier rather than a property that varies with $\ell/m$. The pink curve depicts DQI’s approximation ratio as given by the semicircle law [24], which surpasses the random-assignment bound $r/q$ whenever $\ell/m>0$ and saturates at $\alpha_{\mathrm{DQI}}=1$ for $\ell/m\geq 1-r/q$. In the limit $\ell/m\to 0$, DQI’s performance degrades to exactly $r/q$, matching the worst-case barrier.
We note that on random max-LINSAT(q,r)(q,r) instances with $n$ variables and $m$ constraints, simple algorithms can exceed the $r/q$ threshold. For example Prange’s information set decoding [34] solves an $n\times n$ subsystem exactly and achieves a total expected approximation ratio of $n/m+(1-n/m)\cdot r/q$ for random instances. This does not conflict with Theorem 5: the formula is a heuristic expected-value analysis that assumes the residual mnm-n constraints behave as random evaluations [24, Section 11.3], an assumption that fails on worst-case instances. Moreover, for any fixed $\varepsilon>0$, the PCP-based reduction produces instances in which $n/m$ can be driven below $\varepsilon$, so Prange’s advantage vanishes in exactly the regime where the hardness applies.
This observation is complemented by Theorem 5, which shows that on worst-case max-LINSAT(q,r)(q,r) instances, no polynomial-time algorithm can exceed the $r/q$ threshold (assuming $\mathrm{P}\neq\mathrm{NP}$). While these two statements operate in different regimes – the semicircle law describes DQI’s performance on code-derived instances as a function of the decoding radius, and Theorem 5 applies to adversarially constructed instances with no exploitable structure – they point to the same conclusion from different directions: $r/q$ is the natural worst-case approximation barrier for max-LINSAT, and surpassing it requires both an algorithm capable of exploiting structure and instances that exhibit such structure. The codes for which DQI demonstrates an advantage over known classical algorithms, namely Reed–Solomon codes in OPI [24] and algebraic geometry codes in HOPI [19], satisfy these conditions.
@@
Recent work has further clarified that DQI’s advantage, where it exists, depends critically on problem structure: Ref. [30] shows that for MaxCut instances where DQI achieves any nontrivial asymptotic approximation guarantee, the optimal solution can be found exactly in classical polynomial time. DQI consequently provides no advantage for MaxCut. For unstructured max-$k$-XOR-SAT instances, Ref. [2] proves that DQI is obstructed by the overlap gap property, which is a topological barrier in the near-optimal solution space. Numerical evidence suggests that classical approximate message passing outperforms DQI on such instances. From the perspective of computational complexity, Ref. [29] places DQI in $\mathsf{BPP}^{\mathsf{NP}}$, thereby ruling out immediate quantum advantage via “supremacy”-style hardness arguments as they are used in quantum random sampling [21, 8]. There, it is $\mathsf{P}-$hard to compute individual output probabilities, and an efficient classical sampling algorithm up to a constant error in the total variation distance would lead to a collapse of the polynomial hierarchy to the third level. In contrast, DQI is already placed in a low level of the polynomial hierarchy. The hardness of DQI, if any, arises from locating elements in an exponentially large hidden subset without group structure and not from the hardness of simulating the quantum circuit itself, as here, in contrast to the situation in quantum random sampling, individual probabilities of outcomes can be efficiently computed.
Together with Theorem 5, these results rigorously establish that DQI is not a general-purpose quantum optimizer. Rather, any possible quantum advantage over (known) classical optimizers crucially depends on structures that are present in the instances. Several natural questions remain:
  • Inapproximability for OPI and HOPI: Can inapproximability results be extended to settings studied in OPI [24]? The hard instances produced by our reduction have $n/m=o(1)$ and unstructured constraint matrices, so they do not directly address the Vandermonde-structured instances with constant $n/m$ targeted by DQI. Since maximum-likelihood decoding of Reed–Solomon codes is $\mathsf{NP}$-hard for certain parameters [20], it is natural to ask whether PCP-based methods can be adapted to establish worst-case inapproximability for OPI. The same question extends to the HOPI generalization of Ref. [19].
@@
Our Theorem 5 provides a complexity-theoretic benchmark for max-LINSAT, establishing $r/q$ as the provable worst-case approximation limit and confirming that any improvement must arise from instance structure not present in the hard instances produced by PCP reductions. It is our hope that this result contributes to clarifying the landscape of quantum advantage for optimization problems of practical relevance [17, 23, 1, 6].